# Neuroplasticity in Artificial Intelligence – An Overview and Inspirations on Drop In & Out Learning

### v2

@@ IV-B1 Dropout
Originally, dropout was proposed to prevent overfitting [14], while Baldi et al. [118] provided an alternative perspective, suggesting that dropout can be viewed as a regularisation method applicable to regularised error functions. Wager et al. [119] not only regard dropout as a regularisation technique, but also highlight its connection to AdaGrad [120]. Gal et al. [121] proposed an alternative perspective, interpreting dropout as a Bayesian approximation for modelling uncertainty, as dropout randomly generates sub-networks during training. However, the most conventional form of dropout, which is commonly referenced, is a regularisation technique that randomly sets a proportion of units to zero in NNs during training to prevent overfitting and improve generalisation. The corresponding algorithm is presented in Algorithm 2.
@@
Beyond generalised dropout, specialised dropout strategies have been designed to target specific features in certain neural network architectures. Spatial Dropout [130] was introduced to drop entire feature maps instead of individual neurons, thereby preserving spatial correlations in convolutional neural networks (CNNs). Similarly, RNN Dropout [131] reduces memory consumption by directly dropping recurrent connections in recurrent neural networks (RNNs). A related concept is found in DropConnect [132], where a subset of network weights, rather than individual neurons, is set to zero.
Maxout Dropout [133] was specifically designed for Maxout networks, applying dropout to Maxout units instead of neurons due to the unique structure of this architecture. Likewise, Dropout KAN [134] was developed for Kolmogorov–Arnold Networks (KANs), where learning is based on kernel functions rather than traditional weight-based updates, therefore its dropout operates on post-activations.
@@ IV-C Neuroplasticity: dynamical adjustment of neurons and their connection
The suggested approach further shows some similarities with concepts introduced in the literature: Perwej et al. [8] propose a neuroplasticity-inspired approach in NNs to accelerate training and enhance model performance. Their method involves initially utilising a small network and progressively adding neurons until the AI successfully learns a specific skill, demonstrating neuron addition. Subsequently, they implement a neuron removal process by starting with a large network and gradually eliminating connections while ensuring the model retains its essential skills. Experimental results indicate that pruning two thirds of the connections in a Multilayer Perceptron (MLP) does not considerably degrade performance. Steur et al. [117] propose an additional method that enables Residual Networks and Highway Networks to establish skip connections using Adaptive Nonlinearity Gates (ANGs). Rudroff et al. [116] clearly discuss how the biological system’s ability to balance rapid learning with long-term memory retention can inspire novel AI architectures. They propose a dual-learning architecture: a fast-learning module in the buffer or short-term memory, coupled with offline consolidation replay to the knowledge base or long-term memory. This approach draws inspiration from brain activity during Non-REM (Non-Rapid Eye Movement) and REM sleep, incorporating mechanisms such as SWRs (Sharp Wave Ripples) for reactivation and strengthening, and BARRs (Bidirectional Associative Recall Responses). They present testable hypotheses for both human cognition and neural networks, further bridging AI and neuroscience. These concepts originated from neuroplasticity, which focuses on pruning neurons deemed redundant or non-contributory to the model’s performance. In contrast, approaches inspired by neurogenesis emphasise the activation of new neurons to enhance the model’s overall performances.

### v3

@@ IV-B1 Dropout
Originally, dropout was proposed to prevent overfitting [14], while Baldi et al. [119] provided an alternative perspective, suggesting that dropout can be viewed as a regularisation method applicable to regularised error functions. Wager et al. [120] not only regard dropout as a regularisation technique, but also highlight its connection to AdaGrad [121]. Gal et al. [122] proposed an alternative perspective, interpreting dropout as a Bayesian approximation for modelling uncertainty, as dropout randomly generates sub-networks during training. However, the most conventional form of dropout, which is commonly referenced, is dropout from neurons: a regularisation technique that randomly sets a proportion of units to zero in NNs during training to prevent overfitting and improve generalisation. The corresponding algorithm is presented in Algorithm 2.
@@
Beyond generalised dropout, specialised dropout strategies have been designed to target specific features in certain neural network architectures. Spatial Dropout [131] was introduced to drop entire feature maps instead of individual neurons, thereby preserving spatial correlations in convolutional neural networks (CNNs). Similarly, RNN Dropout [132] reduces memory consumption by directly dropping recurrent connections in recurrent neural networks (RNNs). A related concept is found in DropConnect [133], where a subset of network weights, rather than individual neurons, is set to zero. Maxout Dropout [134] was specifically designed for Maxout networks, applying dropout to Maxout units instead of neurons due to the unique structure of this architecture.
@@
Information Dropout [140] 2018 Dropout based on Information such as feature map
Weight Dropout [141] 2020 Dropout weights directly
@@
Weights dropout
Other than dropout on neurons, previous research has put some efforts on weights dropout. Dropout KAN [135] was developed for Kolmogorov–Arnold Networks (KANs), where learning is based on kernel functions rather than traditional weight-based updates, therefore its dropout operates on post-activations. Weight Dropout, introduced by Sanjar et al. [141], incorporates a dropout mask applied directly to the weights. Their results demonstrate that this method yields improved performance across various tasks.
Information dropout
Beyond focusing solely on dropout in NNs, prior studies have explored broader aspects of learning dynamics in terms of information. One of the earliest discussions appears in [142], which introduces the concept of a critical learning period in deep NNs. This work suggests that, during the initial phase of training, networks experience a surge in information gain, which subsequently diminishes as training progresses. Furthermore, it emphasises the importance of information forgetting as an integral part of the learning process. Building upon this notion, Kleinman et al. [143] argue that multisensory information is progressively integrated within NNs throughout training—a process that exemplifies the model’s capacity for information acquisition. They also maintain that the critical learning period persists in this context.
Therefore, controlling the amount of information at different stages of training plays a crucial role in the training of neural networks. Achille et al. [140] propose a framework for quantifying the complexity of task-relevant information. Based on information-theoretic principles, they introduce Information Dropout, a technique that selectively drops information—such as elements in feature maps—during training. This approach allows the model to automatically adapt to the data and more effectively utilize architectures with limited capacity [144].
@@ IV-C Neuroplasticity: dynamical adjustment of neurons and their connection
The suggested approach further shows some similarities with concepts introduced in the literature: Perwej et al. [8] propose a neuroplasticity-inspired approach in NNs to accelerate training and enhance model performance. Their method involves initially utilising a small network and progressively adding neurons until the AI successfully learns a specific skill, demonstrating neuron addition. Subsequently, they implement a neuron removal process by starting with a large network and gradually eliminating connections while ensuring the model retains its essential skills. Experimental results indicate that pruning two thirds of the connections in a Multilayer Perceptron (MLP) does not considerably degrade performance. Steur et al. [118] propose an additional method that enables Residual Networks and Highway Networks to establish skip connections using Adaptive Nonlinearity Gates (ANGs). Rudroff et al. [117] clearly discuss how the biological system’s ability to balance rapid learning with long-term memory retention can inspire novel AI architectures. They propose a dual-learning architecture: a fast-learning module in the buffer or short-term memory, coupled with offline consolidation replay to the knowledge base or long-term memory. This approach draws inspiration from brain activity during Non-REM (Non-Rapid Eye Movement) and REM sleep, incorporating mechanisms such as SWRs (Sharp Wave Ripples) for reactivation and strengthening, and BARRs (Bidirectional Associative Recall Responses). They present testable hypotheses for both human cognition and neural networks, further bridging AI and neuroscience. Moreover, Zilly et al. [115] observed that weights in neural networks tend to become mutually frozen during sequential task learning, highlighting the limitations of this learning paradigm. To address this issue, they propose reinitializing the connections in the frozen parts of the network, thereby enhancing the plasticity of neural networks and improving their ability to transfer knowledge across tasks. These concepts originated from neuroplasticity, which focuses on pruning neurons deemed redundant or non-contributory to the model’s performance. In contrast, approaches inspired by neurogenesis emphasise the activation of new neurons to enhance the model’s overall performances.